OpenStudy (anonymous):

Dr. Collier summons you over to his table. He wants to demonstrate the graph of a fourth-degree polynomial function, but the batteries in his graphing calculator have run out of juice. Explain to Dr. Collier how to create a rough sketch of a graph of a fourth-degree polynomial function.

OpenStudy (tkhunny):

Well, do it. Don't post a problem with no work shown.

OpenStudy (anonymous):

i don't know how

OpenStudy (anonymous):

that's the question it asked me

OpenStudy (campbell_st):

wow... imagine having to do it by hand..... the old table of values could be used... if you don't know a better method

OpenStudy (anonymous):

this doesn't help me!

OpenStudy (sachintha):

You will have to calculate all the intersection and turning points. :O

OpenStudy (anonymous):

not really all i need to do is explain how to graph it without a calculator

OpenStudy (sachintha):

Well that's what I said.

OpenStudy (tkhunny):

...and what a horrible place this world would be if we did only those thing we already knew how to do. What does y = x^4 look like? What does y = x^4 + 1 look like? What does y = x^4 - 3x look like?

OpenStudy (anonymous):

of what their is no equation given

OpenStudy (campbell_st):

I think you need a little calculus... that may help.... and knowing the difference between P(x) = -x^4........ and P(x) = x^4..... its a question with a lot to consider

OpenStudy (anonymous):

this is stupid because i haven't learn calculus yet and this question isn't even a question lol